OpenStudy (anonymous):

f a sphere’s radius is quadrupled, what is the corresponding change in surface area?

OpenStudy (anonymous):

The surface area is multiplied by 16.... THAT THE ANSWER

OpenStudy (harsimran_hs4):

sa of sphere = 4(pi)r^2 = A so if we make the radius 4 times sa = 4(pi)(4r)^2 = 4(pi)(16r^2) can you figure out now?

OpenStudy (harsimran_hs4):

yes!!